### v1

# Mitigating Metropolitan Carbon Emissions with Dynamic Eco-driving at Scale

Mitigating Metropolitan Carbon Emissions with Dynamic Eco-driving at Scale
Vindula Jayawardana Massachusetts Institute of Technology
{vindula, qua, chickert, edgarrs, cattang, cathywu}@mit.edu
Baptiste Freydt ETH Zürich
bfreydt@student.ethz.ch
Ao Qu Massachusetts Institute of Technology
{vindula, qua, chickert, edgarrs, cattang, cathywu}@mit.edu
Cameron Hickert Massachusetts Institute of Technology
{vindula, qua, chickert, edgarrs, cattang, cathywu}@mit.edu
Edgar Sanchez Massachusetts Institute of Technology
{vindula, qua, chickert, edgarrs, cattang, cathywu}@mit.edu
Catherine Tang Massachusetts Institute of Technology
{vindula, qua, chickert, edgarrs, cattang, cathywu}@mit.edu
Mark Taylor Utah Department of Transportation
{marktaylor, bleonard}@utah.gov
Blaine Leonard Utah Department of Transportation
{marktaylor, bleonard}@utah.gov
Cathy Wu Massachusetts Institute of Technology
{vindula, qua, chickert, edgarrs, cattang, cathywu}@mit.edu
Keywords: Eco-driving, Deep Reinforcement Learning, Semi-autonomous Vehicles, Carbon Emissions, Generalization, Prospective Impact Assessment
1 Abstract
The sheer scale and diversity of transportation make it a formidable sector to decarbonize. Here, we consider an emerging opportunity to reduce carbon emissions: the growing adoption of semi-autonomous vehicles, which can be programmed to mitigate stop-and-go traffic through intelligent speed commands and, thus, reduce emissions. But would such dynamic eco-driving move the needle on climate change? A comprehensive impact analysis has been out of reach due to the vast array of traffic scenarios and the complexity of vehicle emissions. Such an analysis would require careful modeling of many traffic scenarios and solving an eco-driving problem at each one of them - a challenge that has been out of reach for previous studies. We address this challenge with large-scale scenario modeling efforts and by using multi-task deep reinforcement learning with a carefully designed network decomposition strategy. We perform an in-depth prospective impact assessment of dynamic eco-driving at 6,011 signalized intersections across three major US metropolitan cities, simulating a million traffic scenarios. Overall, we find that vehicle trajectories optimized for emissions can cut city-wide intersection carbon emissions by 11-22%, without harming throughput or safety, and with reasonable assumptions, equivalent to the national emissions of Israel and Nigeria, respectively. We find that 10% eco-driving adoption yields 25%-50% of total reduction, and nearly 70% of the benefits come from 20% of intersections, suggesting near-term implementation pathways. However, the composition of this high-impact subset of intersections varies considerably across different adoption levels, with minimal overlap, calling for careful strategic planning for eco-driving deployments. Moreover, the impact of eco-driving, when considered jointly with projections of vehicle electrification, hybrid vehicle adoption, and travel growth, remains significant. More broadly, this work paves the way for large-scale analysis of traffic externalities, such as time, safety, and air quality, and the potential impact of solution strategies. Visual details can be found on the project page https://vindulamj.github.io/eco-drive.
2 Introduction
For interrupted traffic flow dynamics, dynamic eco-driving has attracted significant attention in recent decades Mintsis et al. (2020). Dynamic eco-driving is concerned with reducing carbon emissions and/or fuel consumption of vehicles individually and at the fleet level by strategically modifying driving patterns and behaviors. (Figure 1). Owing to advances in vehicle-to-infrastructure (V2I) communication technologies and the standardization of the signal phase and timing (SPaT) messages, these technologies have facilitated the research of dynamic eco-driving, especially near signalized intersections, where traffic flow is persistently interrupted Mintsis et al. (2020). Based on prior studies, we crudely estimate that persistent yet unproductive idling and excess accelerations at signalized intersections could contribute up to 14.6% of land transportation CO2 emissions in the United States (US) (refer to Appendix 10.1). Highlighting its significance, this is comparable to half of the well-scrutinized US airline sector emissions Lee et al. (2009). Reducing emissions at signalized intersections thus emerges as a promising avenue for achieving climate change mitigation goals.
Figure 1: A schematic illustration of eco-driving at signalized intersections. Controlled vehicles are driven in a way that reduces their carbon emissions. Through car following dynamics and implicitly forming local cooperative teams, controlled vehicle behaviors can affect human-driven vehicle behaviors, leading to a reduction in carbon emissions in human-driven vehicles as well.
However, despite nearly four decades of research, large-scale implementation of dynamic eco-driving is yet to be seen. A potential reason and a notable observation is that existing studies vary widely in assumptions and, thus, reported outcomes. In particular, previous studies report a significant variation in emission benefits ranging from 2% to 56% Mintsis et al. (2020). Furthermore, our observation is that a majority of studies are small-scale (e.g., one or two intersections or a few intersections in a route) and mainly focus on technical questions that concern how to eco-drive Yang et al. (2016, 2020); Sun et al. (2020). But the fundamental socio-technical question of whether to eco-drive, which requires going beyond a few intersections-based analyses, remains unanswered Barkenbus (2010). Our aim of this work is to contribute a large-scale prospective analysis of eco-driving at signalized intersections, highlighting city-scale emission benefits and uncovering insights that may influence widespread adoption—factors that are often elusive in smaller-scale studies. Such an analysis provides a scientific basis for informed decision-making, strategic planning of infrastructure, and technology development Barkenbus (2010).
However, conducting a large-scale analysis of eco-driving presents significant challenges, a complexity that has largely eluded in previous studies. Diverse eco-driving scenarios stemming from factors such as intersection topologies, vehicle-related factors (e.g., fuel type, age, model, etc.), weather conditions, and seasonal changes must be considered to ensure that the prospective eco-driving impact assessments are representative Falkenauer (1998); Jayawardana et al. (2022). It not only ensures that insights are generalizable but also aligns with industry best practices for impact assessments iee (2022); Ross and Guhathakurta (2017); Duinker and Greig (2007). While field studies are considered a gold standard, as they reflect real-world conditions, it is prohibitively expensive to conduct traffic field tests at scale Gloudemans et al. (2023). Simulation-based analyses offer greater scalability, and uncertainties stemming from the sim-to-real gap can be addressed Vinitsky et al. (2022); Jang et al. (2019).
Simulation-based analysis consists of three key steps. First, each scenario is modeled within a high-fidelity agent-based traffic simulator, replicating real-world intersections and calibrating for real-world conditions. This step is termed scenario modeling. Second, a set of eco-driving strategies is devised for each scenario to control the vehicles based on a given eco-driving adoption level, referred to as eco-driving control. Lastly, the impact of eco-driving using the devised strategies on the modeled scenarios is assessed and compared against naturalistic human-like driving as a baseline. It is termed impact assessment.
We identified 33 influential factors for eco-driving based on previous studies and expert opinion. An eco-driving scenario manifests as a combination of these factor valuations. Each factor can take multiple valuations, creating a vast scenario space of at least $2^{33}$ scenarios. A comprehensive eco-driving analysis thus suffers from the curse of dimensionality Donoho et al. (2000). The high-dimensional scenario space grows exponentially with each additional factor, amplifying computational complexity in scenario modeling and eco-driving control. Furthermore, in eco-driving control, each scenario involves solving an optimal eco-driving control problem, making a comprehensive large-scale analysis challenging due to the need to solve a vast number of them—a scale far out of reach of previous studies.
This article addresses these challenges to comprehensively analyze eco-driving at signalized intersections at scale. We consider semi-autonomous vehicles, such as vehicles equipped with advanced driver assistance systems (ADAS), which we can be leveraged to intelligently manage stop-and-go traffic, effectively reducing carbon emissions. To navigate the complexity of this undertaking, we take a data-driven approach to scenario modeling and model the scenario space with representative scenarios instead of exhausting all scenarios with inputs from domain experts. We then optimize eco-driving control across this––still significantly large––spectrum of representative scenarios by employing model-free multi-task deep reinforcement learning and zero-shot transfer learning instead of often used model-based methods. To make the optimization of eco-driving control tractable, we employ a carefully designed network decomposition strategy. This approach focuses on optimizing eco-driving at each intersection individually, instead of network scope optimizations, which is a significantly hard problem and intractable even with advances in modern computing and thus remains an open optimization challenge Qadri et al. (2020). While network decomposition strategies are commonly used in eco-driving research, many studies fail to analyze the impact of eco-driving optimized on individual intersections when deployed at the network scale. For example, if the throughput is increased due to eco-driving at an optimized intersection, that may lead to a queue spillback event in the following intersection at the network level. This oversight can render the previous analyses less reliable, potentially leading to network-level traffic overflows when such strategies are implemented. In contrast, we ensure that such impacts are accounted for in the assessment to ensure that the benefits observed in our analysis are consistent and reliable at the network level.
Analyzing nearly one million traffic scenarios stemming at 6,110 intersections across three major high-emissions US metropolitan cities Gately et al. (2015)- Atlanta, Los Angeles, and San Francisco- our analysis reveals insights that were elusive at small scale analyses and strengthen the existing knowledge with high confidence. We find that implementing eco-driving could reduce US intersection CO2 emissions by 11-22%, which under reasonable extrapolating assumptions is equivalent to the national carbon emissions of Israel and Nigeria (Refer to Appendix 10.2). Such an implementation does not affect intersection throughput and could even assist in increasing throughput due to smoothed traffic flow. Furthermore, we observe that eco-driving is as safe as conventional driving, measured by safety surrogate measures Wang et al. (2021a). With just 10% of eco-driving adoption, 25%-50% of the total emission benefits (benefits if the adoption is 100%) can be achieved, underscoring its significant near-term potential as a climate change mitigation strategy. However, when eco-driving adoption increases, the impact level of each factor affecting eco-driving benefits changes. For instance, while longer green signal times may negatively affect benefits at low adoption levels, they can positively contribute at higher adoption rates. This dynamic relationship necessitates frequent adjustments in traffic engineering practices and a reevaluation of long-term infrastructure investments at intersections. Furthermore, approximately 70% of the total benefits are concentrated in just 20% of intersections, even at lower adoption levels. However, the composition of this high-impact subset of intersections varies considerably across different adoption levels, with minimal overlap. This finding underscores the need for a more nuanced and adaptive approach to eco-driving deployment, emphasizing the importance of careful planning and strategic implementation to maximize benefits as adoption rates evolve over time.
Diverging from the sole technical focus often found in eco-driving research, this article adopts a socio-technical perspective, seeking to contextualize eco-driving within the framework of societal priorities. In a broader context, this work paves the way for scalable analyses of traffic externalities and solution strategies—a critical opportunity in light of the expanding adoption of roadway automation. Below, we summarize our main contributions,
  • We present, to the best of our knowledge, the first prospective impact assessment of metropolitan city-scale eco-driving at signalized intersections, considering three main cities in the United States - Atlanta, Los Angeles, and San Francisco.
  • We perform data-driven modeling of one million traffic scenarios at 6,110 intersections in the three cities in an agent-based microscopic traffic simulator.
  • We devise generalizable eco-driving policies for vehicles using multi-task deep reinforcement learning and zero-shot transfer learning across these scenarios with a carefully defined network decomposition strategy.
  • We conduct a comprehensive assessment of eco-driving at signalized intersections using the modeled scenarios and devised control policies and show insights that are elusive in small-scale analyses.
3 Related Work
In the following sections, we present a summary of related work by categorizing them into four sub-sections that are closely related to our work.
3.1 Eco-driving at signalized intersections
Over the past decade, numerous studies have investigated the challenge of dynamic eco-driving at signalized intersections. While they often study different traffic settings, Mintsis et al. Mintsis et al. (2020) show an average eco-driving benefit of 17%, with a standard deviation of 11% by considering 21 past studies. While this highlights the potential of eco-driving, these studies often consider only a few factors that are known to have an impact on eco-driving benefits, making the shown benefits scenario/traffic setting specific. Based on the literature, we identify nearly 33 factors that affect eco-driving benefits. We refer the reader to Appendix 10.3 for more details on these factors and their known impact.
Moreover, most eco-driving studies focus on devising better eco-driving strategies (e.g., how to eco-drive). Such studies often take three forms: 1) analytical approaches, 2) simulation approaches, and 3) field experiments Mintsis et al. (2020). Analytical approaches often focus on analyzing the eco-driving problem from a theoretical point of view by analyzing properties such as the stability of the system under certain eco-driving strategies Coppola et al. (2022). On the other hand, simulation-based approaches are used to test the potential of new eco-driving strategies and also to conduct impact assessments. However, one of the major limitations of such work is the lack of consideration of many factors that affect eco-driving benefitsBarkenbus (2010); Mintsis et al. (2020); Jayawardana and Wu (2022); Meng and Cassandras (2020). Furthermore, for simplicity and tractability of eco-driving control optimization scope, most of these studies analyze a few intersections under simplified traffic conditions (e.g., uniform inflow rates) Yang et al. (2016, 2020); Sun et al. (2020). A few studies that focus on multiple intersections often solely focus on the generalization of eco-driving strategies and rely on synthetically generated traffic scenarios while not considering the impact assessment requirements Jayawardana et al. (2024). Finally, field studies are also conducted to demonstrate the benefits of eco-driving and to validate the applicability of eco-driving strategies Hao et al. (2018); Stahlmann et al. (2018); Mintsis et al. (2017). However, these studies are often confined to small regions (e.g., a few intersections) due to the logistical difficulties of conducting such experiments at scale Mintsis et al. (2020).
3.2 Traffic intersection scenario modeling
Numerous studies have focused on modeling intersections for various purposes, such as traffic signal control Ault and Sharon (2021) and autonomous driving Leurent (2018). However, they often remain limited to specific scenarios Wang et al. (2021b); Ault and Sharon (2021), primarily due to a focus on proof of concepts rather than comprehensive assessments of proposed methods. Additionally, modeling multiple intersections can be prohibitively expensive, requiring data fusion from various sources and handling missing data.
However, our requirements for modeling intersection scenarios are objectively and strategically different from previous work, necessitating us to develop our own datasets. First, we are neither presenting an eco-driving proof of concept nor attempting to answer how to do eco-driving. Our goal is to analyze the impact of eco-driving to answer the question of whether to eco-drive. Such questions can not be answered by analyzing merely a few intersection scenarios. Unless we have a representative set of intersection scenarios, merely relying on a select intersection could lead to evaluation overfitting where the insights extracted from a few scenarios do not generalize to the class of scenarios Jayawardana et al. (2022); Whiteson et al. (2011).
Furthermore, we must avoid overfitting to a specific region in our impact assessment Gately et al. (2015). While there are popular intersection datasets available for Manhattan Island in New York Wei et al. (2019), using them could lead to regional overfitting. Manhattan Island exhibits distinct spatial and temporal traffic dynamics due to its block-oriented road network. Insights drawn from an impact assessment solely based on these intersections would not be effective in generalizing to other cities or regions.
Another option is to use synthetic datasets, which are easier to generate Li et al. (2022). However, impact assessments based on these datasets are not particularly meaningful as they may not reflect the real-world intersection distributions and may contain oversimplified scenarios Feng et al. (2023).
Thus, the fundamental requirement for modeling intersection scenarios lies in their diversity, encompassing variations at both the intersection and regional levels and making sure they are drawn from the real-world distribution. However, it is crucial to exercise caution when creating datasets, ensuring that the fused data originate from the same distribution. For instance, the IntD dataset Bock et al. (2020) containing a set of intersections in Germany is popular within the research community, but combining it with U.S.-specific data like vehicle inflows may result in a distribution mismatch.
Thus, to ensure the accuracy and validity of the modeling process, we must meticulously consider the sources and characteristics of the data we fuse, striving to maintain consistency within the datasets to avoid any potential biases or inaccuracies and be explicit about the assumptions we make. In this work, we follow these steps to carefully approach the scenario modeling while incorporating feedback from domain experts, as will be explained later.
3.3 Methods for eco-driving control
Earlier studies on eco-driving have predominantly explored five control strategies for controlling vehicles. They include 1) dynamic programming, 2) heuristics, 3) trajectory optimizations, 4) model-based reinforcement learning (model-based RL), and 5) model-free reinforcement learning (model-free RL). Each of these approaches exhibits distinct advantages and drawbacks. In Table 1, we offer a comparative analysis of these methods, focusing on five prevalent shortcomings.
Method Sub opt- Rely on Model-based High comp- Small
imality estimations assumptions uting load scale
Dynamic programming[1,2] x x x x
Heuristics[3] x x
Trajectory optimization[4,5] x x * *
Model-based RL[6,7] * * x *
Model-free RL[8-10] * *
Table 1: A comparison of the shortcomings of five known methodologies for eco-driving control policy synthesis. The ’✓’ indicates that the method does not demonstrate the shortcoming, an ’x’ indicates that the method demonstrates the shortcoming, and a ’*’ indicates that the shortcoming may or may not present. [1]Maamria et al. (2016), [2]Doan et al. (2018), [3]Katsaros et al. (2011), [4]Kamal et al. (2010), [5]Wang et al. (2019), [6]Lee et al. (2020), [7]Zhu et al. (2022), [8]Jayawardana and Wu (2022), [9]Wegener et al. (2021), [10]Bai et al. (2022)
  • Sub optimality: Refers to the method’s ability to produce optimal results.
  • Rely on estimations: Refers to having access to additional information, like intersection queue lengths derived from fundamental diagrams during the optimization stage. This shortcoming encompasses both data acquisition and communication of those data-related estimations.
  • Model-based assumptions: Involves relying on a potentially complex model of vehicle dynamics, especially challenging when dealing with inter-vehicle dynamics and lane changes. Simplified models can lead to sub-optimal solutions.
  • High computational load: Indicates the run-time computational burden needed for obtaining solutions. This is crucial, considering that vehicles may not have powerful onboard computers.
  • Small scale: Refers to the optimization problem’s scale, e.g., one intersection or two. It does not address deployment scale directly, although some synergies may exist.
As can be seen from Table 1, model-free deep reinforcement learning appears as a favorable choice among the alternatives. Thus, in this work, we leverage model-free deep reinforcement learning to devise eco-driving strategies.
3.4 Reinforcement learning for vehicle-based traffic control
Contrary to conventional traffic control mechanisms such as traffic signals and stop signs, vehicle-based traffic control is increasingly being explored in recent research. Such vehicle-based control is referred to as Lagrangian traffic control and involves controlling a fleet of vehicles to achieve certain fleet-level (including both controlled and uncontrolled vehicles) objectives such as congestion management Vinitsky et al. (2018), reducing emissions Jayawardana and Wu (2022), and fostering smoother traffic flow Wu et al. (2021); Yan et al. (2022). While various studies have presented proofs-of-concept, they still fall short of accurately modeling real-world scenarios Yan et al. (2022); Wu et al. (2017); Di and Shi (2021) or go beyond toy traffic problems Yan et al. (2022).
Existing research often relies on simplified traffic models, such as toy traffic blocks Yan et al. (2022), to assess the benefits of Lagrangian traffic control methods. However, none of these studies explore the potential of large-scale implementations that could offer meaningful insights for practical decision-making. While toy tasks are valuable for gaining preliminary insights into a method, they might not resonate strongly with policymakers and stakeholders due to the lack of rigorous analysis of the benefits at a scale that holds significance Esser and Rickert (2000).
In order to create more impactful outcomes that can drive policy changes and engage stakeholders, future research should focus on conducting comprehensive analyses that consider real-world complexities and practical scenarios. By doing so, the potential benefits of Lagrangian traffic control can be better understood, and its implementation may gain the attention and support needed to truly transform transportation systems, and make a substantial positive impact on traffic congestion and emission reduction, and improve overall traffic flow dynamics. In this work, we take a step towards this direction by analyzing the impact of Lagrangian eco-driving at signalized intersections in the metropolitan city scale.
4 Preliminaries
4.1 Reinforcement Learning
In reinforcement learning, an agent learns a control policy by interacting with its environment, typically modeled as a Markov Decision Process (MDP) denoted as $M=\left\langle\mathcal{S},\mathcal{A},p,r,\rho,\gamma\right\rangle$. Here, $\mathcal{S}$ represents the set of states, $\mathcal{A}$ denotes the possible actions, $p(s_{t+1}|s_{t},a_{t})$ denotes the transition probability from the current state $s_{t}$ to the next state $s_{t+1}$ upon taking action $a_{t}$, the reward received for action $a_{t}$ at state $s_{t}$ is $r(s_{t},a_{t})\in\mathbb{R}$, a distribution over the initial states is $\rho$, and $\gamma\in[0,1]$ is a discounting factor that balances immediate and future rewards.
Given the MDP, we seek to find an optimal policy $\pi^{*}:\mathcal{S}\rightarrow\mathcal{A}$ over the horizon $H$ that maximizes the expected cumulative discounted reward over the MDP.
$\pi^{*}(s)=\operatorname*{argmax}_{\pi}\mathop{\mathbb{E}}\left[\sum_{t=0}^{H}{\gamma^{t}r(s_{t},a_{t})|s_{0},\pi}\right]$ (1)
4.2 Contextual Reinforcement Learning
A Contextual Markov Decision Process (cMDP) is an extension to the definition of an MDP introduced in Section 4.1 by incorporating contextual information. The context parameterizes environmental variations, like changes in speed limits or lane lengths at different intersections in eco-driving. A cMDP is denoted as $\mathcal{M}=\left\langle\mathcal{S},\mathcal{A},\mathcal{C},p_{c},r_{c},\rho_{c},\gamma\right\rangle$. Different from MDPs, a context space $\mathcal{C}$ is introduced, but the action space A and state space S stay the same. The transition dynamics $\textit{p}_{c}$, rewards $\textit{r}_{c}$, and initial state distribution $\rho_{c}$ are conditioned on the context $c\in\mathcal{C}$. In other words, a cMDP $\mathcal{M}$ is a set of MDPs, each defined based on a context, and =$\mathcal{M}=\{M_{c}\}_{c\sim\mathcal{C}}$ where $M_{c}$ is an MDP.
Solving a cMDP requires solving an algorithmic generalization problem within the task. This can be achieved by finding a policy that performs well in the cMDP overall, as denoted by Equation 2.
$\pi^{*}(s)=\operatorname*{argmax}_{\pi}\mathop{\mathbb{E}}\left[\sum_{c\in\mathcal{C}}\sum_{t=0}^{H}{\gamma^{t}r_{c}(s_{t},a_{t})|s_{0}^{c},\pi}\right]$ (2)
4.3 Multi-task Learning
In multi-task reinforcement learning, we extend the single-MDP (single task) reinforcement learning in Section 4.1 to multiple MDPs (multiple tasks). Accordingly, our objective in finding optimal policy thus becomes,
$\pi^{*}(s)=\operatorname*{argmax}_{\pi}\mathop{\mathbb{E}}\left[\sum_{\tau\in\mathcal{T}}\sum_{t=0}^{H}{\gamma^{t}r_{\tau}(s_{t},a_{t})|s_{0}^{\tau},\pi}\right]$ (3)
where $\mathcal{T}$ is the set of MDPs (tasks). When this set of MDPs is defined based on a context, multi-task learning solves a cMDP. Also, note that we seek a unified policy in multi-task reinforcement learning that is performant over all MDPs (tasks).
4.4 Zero-shot Transfer Learning
Zero-shot transfer learning extends the concept of learning across multiple tasks, as in cMDPs, where the agent must perform well on new, unseen tasks without additional training. The goal is to leverage knowledge from a set of source tasks $\mathcal{T}_{\text{source}}$, which the agent has experienced during training, to generalize effectively to a set of target tasks $\mathcal{T}_{\text{target}}$. These target tasks are typically drawn from a distribution that may encompass tasks not encountered during training. In particular, a policy $\pi_{source}$ trained with the $\mathcal{T}_{\text{source}}$ is evaluated on the target tasks $\mathcal{T}_{\text{target}}$ without further training on the target tasks. Such a transfer learning strategy reduces the training overhead and can be better than training a policy from scratch on the target tasks Kirk et al. .
5 Problem Formulation
5.1 Regional eco-driving problem
Optimizing eco-driving at the traffic network scale can yield the highest benefits, but it is a significantly hard problem and intractable even with advances in modern computing and thus remains an open optimization challenge Qadri et al. (2020). Therefore, many studies adopt a network decomposition from the full network to the intersection level and optimize vehicle trajectories at individual intersections. While this has been a standard practice, it has to be performed carefully as isolated optimization of intersection-level behaviors can cause unexpected adverse events (e.g., intersection queue spillback from one intersection to another) at the network level. In this work, we adopt the same network decomposition strategy and take careful precautions to isolate the effects of the changed vehicle behaviors on each individual intersection.
In light of this, consider a region comprised of traffic scenario variations $\Phi$ where $|\Phi|$ is very large (e.g., millions of individual signalized intersection scenarios). With the network decomposition strategy, traffic scenario variations are stemming at the intersection level (e.g., changes in traffic flow, temperature, etc.). Let $\pi^{e}\coloneqq(\pi^{e}_{\phi})_{\phi\in\Phi}$ denote the set of eco-driving control laws. Similarly, let $\pi^{b}\coloneqq(\pi^{b}_{\phi})_{\phi\in\Phi}$ denote the baseline (status quo driving). Let $f(\pi,\phi)$ denote a function that captures the CO2 emission for the scenario $\phi$ under the control law $\pi$. We then define the regional eco-driving effectiveness as,
$E_{\Phi}(\pi^{e},\pi^{b})\coloneqq 1-\frac{\operatorname*{\mathbb{E}}_{\phi\in\Phi}[f(\pi^{e}_{\phi},\phi)]}{\operatorname*{\mathbb{E}}_{\phi\in\Phi}[f(\pi^{b}_{\phi},\phi)]}$ (4)
We thus seek to solve the regional eco-driving problem for the control laws $\pi^{e}$ such that,
Find
$\pi^{e}=\arg\max_{\pi}E_{\Phi}(\pi,\pi^{b})$ (5)
subject to,
$n_{\phi}^{e}\geq n_{\phi}^{b}\;\;\;\;\forall\;\phi\in\Phi$ (6)
$m_{\phi}^{e}\leq\tau\;\;\;\;\forall\;\phi\in\Phi$ (7)
$w_{\phi}^{e}\leq w_{\phi}^{b}\;\;\;\;\forall\;\phi\in\Phi$ (8)
Here, the $n_{\phi}^{e}$ and $n_{\phi}^{b}$ denote the throughput of the intersection where the eco-driving is assessed in the scenario $\phi$ under the control laws of $\pi^{e}_{\phi}$ and $\pi^{b}_{\phi}$, respectively. Similarly, $m_{\phi}^{e}$ denotes the average observed queue length to lane length ratio of the immediately following intersection to the intersection where the eco-driving is assessed, and $\tau$ is a hyperparameter. Finally, $w_{\phi}^{e}$ and $w_{\phi}^{b}$ denote the average waiting time per vehicle in the immediate preceding intersection to the intersection where the eco-driving is assessed under control laws of $\pi^{e}_{\phi}$ and $\pi^{b}_{\phi}$, respectively.
The Equations 6, 7, and 8 ensure the network decomposition strategy would not induce traffic events that would adversely affect the traffic flow at the network level. In particular, Equations 6 ensures that by eco-driving, we do not reduce the intersection throughout, as a reduction in throughput could create traffic queues that would adversely affect the network traffic flow. Equations 7 ensure that if there is an increase in the throughput at an intersection with eco-driving, it will not lead to a saturated incoming approach in the immediately following intersection, which ensures that the immediately following intersection has the capacity to absorb the increased throughput. Finally, Equation 8 ensures that if the eco-driving affects the inflow of vehicles to the intersection being assessed, it would not be propagated back to the intersection preceding it. These constraints collectively ensure that changes in the traffic flow due to eco-driving are either localized to the assessed intersection or implemented with high confidence that they will not adversely affect overall network traffic flow. While this conservative approach may yield a lower bound of potential benefits, it provides a reliable assessment of the impact of eco-driving considering the network scale traffic flow dynamics.
5.2 Eco-driving control problem at an intersection
Finding the policy $\pi^{e}$ as described in Section 5.1, leads to solving an eco-driving control problem for every traffic scenario. Here, we offer a high-level overview of this eco-driving control problem for a generic scenario. It is important to note that every scenario demands a customized formulation to address its unique instantiations of the eco-driving factors.
The main objective of the optimal eco-driving control problem for a given scenario at a signalized intersection is to minimize the emissions of a fleet of vehicles (both controlled semi-autonomous vehicles and human-driven vehicles). However, this should be done with minimal impact on the travel times of vehicles. Given a vehicle dynamics model $x_{i}$ for guided semi-autonomous vehicle $i$ and $y_{j}$ for human-driven vehicle $j$ and the traffic signal timing plan $p$ of the signalized intersection being assessed, we formulate the optimal-control problem in discrete time-space as follows,
$\min J=\sum_{i=1}^{n}\sum_{t=0}^{T_{i}}E\left(a_{i}(t),v_{i}(t)\right)+T_{i}$ (9)
such that for every vehicle $i$,
$a_{i}(t)=\begin{cases}x_{i}(s_{i}(t),\dot{s}_{i}(t),v_{i}(t),p)&\\ \quad\quad\quad\quad\quad\quad\text{if vehicle $i$ is controlled to eco-drive}\\ y_{i}(s_{i}(t),\dot{s}_{i}(t),v_{i}(t),p)&\\ \quad\quad\quad\quad\quad\quad\text{otherwise}\end{cases}$ (10)
$\sum_{0}^{T_{i}}v_{i}(t)\Delta t=d$ (11)
$s_{min}\leq s_{i}(t)\leq s_{max}\quad\forall t\in[0,T_{i}]$ (12)
$v_{min}\leq v_{i}(t)\leq v_{max}\quad\forall t\in[0,T_{i}]$ (13)
$a_{min}\leq a_{i}(t)\leq a_{max}\quad\forall t\in[0,T_{i}]$ (14)
Here, $n$ represents the total number of vehicles in the fleet, $a_{i}(t)$, $v_{i}(t)$, $s_{i}(t)$, and $\dot{s}(t)$ denote the acceleration, velocity, headway and relative velocity of vehicle $i$ at time $t$, $E(\cdot)$ denotes the emission function that measures instantaneous emission of vehicle $i$ at time $t$, $T_{i}$ denotes the travel time of vehicle $i$, and $d$ represents the total required travel distance, which includes both arrival to and departure from the intersection. Furthermore, $\Delta t$ defines the discretized time step usually in $[0.5,1]$.
The minimum and maximum headway values are denoted as $s_{min}$ and $s_{max}$, while $v_{min}$, $v_{max}$ and $a_{min}$, $a_{max}$ indicate the same for velocities and accelerations. These variables and constraints ensure safety, connectivity via V2V communication (if applicable), and passenger comfort.
We use IDM Treiber et al. (2000) as the dynamics model for human drivers. Thus, $y_{j}$ entails the ordinary differential equations defined by the calibrated parameterization of IDM for human-driven vehicle $j$. The $x_{i}$ for controlled vehicle $i$ is learned using multi-task deep reinforcement learning.
Figure 2: An overview of all steps included in assessing the impact of eco-driving. The high-level steps include scenario modeling, eco-driving control, and impact assessment. In scenario modeling, a set $\Phi$ of traffic scenarios is identified based on the factors that affect eco-driving benefits. These scenarios are pruned to select a subset $\Phi^{\prime}$ of representative scenarios. Another subset $\phi_{T}$ is selected for training eco-driving policies. By using the subset $\Phi_{T}$, we then train a set of eco-driving policies $\Pi$ using multi-task deep reinforcement learning. These policies are then used for zero-shot transfer learning-based policy selection to select the best policy for each scenario in $\Phi^{\prime}$. This final scenario to policy assignment is denoted by $\Pi^{e}$. Finally, the impact assessment is performed using $\Pi^{e}$ on $\Phi^{\prime}$.
6 Methodology
Figure 3: A visual illustration of eco-driving factors and how they affect emissions modeling. The legend on the top indicates how each column is connected. Emission cause: Factors stem from different emission causes. A single factor can manifest due to multiple causes, but we illustrate a known predominant cause for simplicity. Factor: there are around 33 major factors affecting emissions of vehicles. Factor valuations: A list of factor values we consider as each factor can take multiple values. For simplicity of the figure, we omit factor values for continuous variable factors. Modeling choice: Indicates what knowledge or data sources inform each factor/factor value. Methodology: Indicates how we translate the modeling choice into our simulation-based analysis. We leverage three levels of modeling: traffic-level, vehicle-level, and behavior-level. Each modeling level is captured by one modeling methodology.
In the following subsections, we detail our approach to assessing the impact of eco-driving at signalized intersections by describing our approach to scenario modeling, solving eco-driving control, and how we perform impact assessment. A visual illustration of the overall methodology and how each of the three steps are connected is given in Figure 2.
6.1 Scenario Modeling
Scenario modeling involves modeling each eco-driving scenario within a high-fidelity agent-based traffic simulator, replicating intersections as digital intersections, and calibrating for real-world conditions. Naive scenario modeling, which exhaustively models all possible scenarios, lacks practicality due to the inclusion of unrealistic scenarios (e.g., snow in Phoenix, Arizona). On the other hand, exhaustive yet realistic modeling offers superfluous detail that is redundant in decision-making (e.g., temperature impact on eco-driving at a fine-grained scale). Our approach seeks a representative subset of scenarios ($\Phi^{{}^{\prime}}\subset\Phi$), ensuring realism, computational feasibility, and comprehensive coverage, as shown in Figure 4.
As illustrated in Figure 3, related literature and expert opinions identify 33 major factors that impact eco-driving benefit levels. We omit the details of each factor here and refer the reader to Section 10.3 in the Appendix for further details on each of these 33 factors and their known effects. Given the diverse spectrum of valuations each factor can take (e.g., seasons spanning summer, fall, winter, or spring), an eco-driving scenario manifests as a combination derived from the potential values across all factors. This results in an exponentially expanding scenario tree, as in Figure 5.
Figure 4: Diverse scenario modeling choices and their relative comparison in terms of how many scenarios each modeling choice would require to model (not drawn to a scale).
The scenario tree yields an exponential number of scenarios (at least $2^{33}$ scenarios). Our first goal is to simplify this scenario tree, balancing representativeness and the tractability of analysis. For this purpose, we prune the potential values each factor can take based on their significance, as known in the literature and based on the authors’ assessment of how much additional information we would obtain if we were to assess a given scenario in the presence of other already selected scenarios. For instance, we take average seasonal temperatures per city instead of every temperature variation, which drastically reduces the tree branching involving the temperature factor. This is motivated by the observation that assessing all temperature variations will result in redundant information and little to no additional knowledge beyond analyzing a selected few temperatures. This curation involves expert consultation and extensive literature review, resulting in a subset $\Phi^{{}^{\prime}}\subset\Phi$ of one million scenarios ($2^{20}$), representing a scenario space reduction by orders of magnitude. Due to the scale of this curation, we omit details here and refer the reader to Appendix Tables 4 5, 6, and 7 and Section 10.4 for detailed curated modeling of each factor and Appendix 10.6 for a high-level breakdown of the one million scenarios. Below, we provide a high-level overview of our approach in scenario modeling.
Figure 5: Eco-driving scenarios tree using eco-driving factors. Each path from the root to a leaf node indicates a scenario, whereas green paths indicate representative scenarios.
For any given scenario, we capture the emissions effects of these factors through three interrelated levels of modeling: 1) traffic-level, 2) vehicle-level, and 3) behavior-level as shown in Figure 3. All simulation scenario modeling is performed in the open-source agent-based SUMO microscopic traffic simulator Lopez et al. (2018).
Traffic-level modeling involves factors owing to the traffic scenario. These factors are modeled using traffic microsimulation. This includes using real-world intersection data to model intersection layouts, speed limits and vehicle inflows, vehicle age, vehicle type, fuel type distributions, etc. For human driver behaviors, we calibrate car-following models using real-world driving data. Since traffic flow at intersections is interconnected, with outflows from one affecting another, we model neighboring intersections where eco-driving is not active as ghost intersections, while the intersection where eco-driving is active is termed the control active intersection as illustrated in Figure 6. We consider optimal fixed-time traffic signal plans and assume pedestrian and cyclist compliance with road rules.
In vehicle-level modeling, we aim to capture the factors owing to the operating characteristics of individual vehicle engines. These factors are modeled using vehicle emission models. This approach allows us to account for factors such as temperature, humidity, seasonality, and weather conditions. The industry-standard Motor Vehicle Emission Simulator (MOVES) by the US Environmental Protection Agency epa (a) provides accurate exhaust emissions for this purpose. Yet, it is impractical for our large-scale analysis, where we must assess millions of scenarios, given the MOVES takes about three minutes per scenario. To overcome this, we create a suite of efficient surrogate neural network emission models using MOVES as the data source Sanchez et al. (2022, 2023). Leveraging MOVES also allows us to capture the cumulative average effects of vehicle maintenance, vehicle model, vehicle characteristics, road types, aerodynamic drag, air conditioning, and vehicle load.
Last, we focus on behavior-level modeling, which accounts for the factors owing to altered vehicle behaviors due to eco-driving optimization. These factors are modeled using deep reinforcement learning and rule-based control modeling. We consider behaviors with a substantial impact on emissions, such as idling, speeding, acceleration, deceleration, pulse-and-glide, and lane changes Xu et al. (2021). Behavior-level modeling across diverse traffic scenarios is a core challenge addressed in this work, as discussed next.
6.2 Eco-driving Control
Figure 6: The control active intersection (where the vehicles are controlled for eco-driving), and ghost intersections (where the vehicles are not controlled for eco-driving). Each incoming approach is preceded by a ghost approach.
Eco-driving control involves devising a set of eco-driving strategies for the representative scenario set $\Phi^{{}^{\prime}}$. Eco-driving control for one million scenarios is a challenging task. Traditional methods like dynamic programming are computationally limiting Doan et al. (2018), while heuristics are often suboptimal Katsaros et al. (2011). Trajectory optimizations are limited by the complexity of required vehicle dynamics models, particularly due to inter-vehicle dynamics and environmental variability introduced by many factors we consider Yang et al. (2016).
Meanwhile, recent studies have shown the effectiveness of deep reinforcement learning in complex control tasks Wurman et al. (2022); Silver et al. (2017); Bellemare et al. (2020). Neural vehicle control using deep reinforcement learning Wu et al. (2021); Yan et al. (2022); Jayawardana and Wu (2022), which can handle diverse traffic scenarios without manual fine-tuning, is also gaining popularity. However, training such a policy for each scenario is impractical due to the number of scenarios. Training a unified policy is complex due to multifaceted factors and the induced complex expectation Zhang and Yang (2021). Learning both discrete lane changes and continuous longitudinal accelerations presents optimization challenges, demanding diverse exploration strategies and distinct neural network representations.
Thus, we introduce a custom training method involving three key steps: 1) similar scenario partitioning, 2) within partition multi-task deep reinforcement learning (MT-DRL), and 3) rule-based lane changes and a hierarchical policy selection strategy to accommodate lane changes. Below, we detail each of these steps in detail.
Similar scenario partitioning: In this step, our aim is to partition scenarios by their similarities, allowing us to train separate MT-DRL policies for each partition. This strategy is inspired by the mixture of expert architecture widely used in real-world systems with generalization requirements Masoudnia and Ebrahimpour (2014). It can reduce scenario diversity within a partition and help reduce training complexity. While each scenario is defined by many factors, we only utilize intersection geometry and traffic flow-specific features as partitioning features because they strongly influence policy behaviors.
However, since signalized intersections lack rotational invariance, comparing them for similarity is challenging. Therefore, we break down the intersection environment into non-conflicting pairs of incoming and outgoing approaches. This has added advantages. It accelerates training by simulating fewer vehicles and improves training stability by alleviating credit assignment challenges Zhou et al. (2020). Further, it has no impact on modeling realism due to the decoupling nature of traffic signals (e.g., northbound traffic and eastbound traffic do not collide). These selected pairs, along with the ghost approaches connected to them, thus become our training environments and are illustrated in Figure 7.
Figure 7: A training environment is defined by a pair of incoming and outgoing approaches with the corresponding ghost approaches. In each training environment, a realistic traffic inflow to the control active intersection is generated using its incoming approaches and ghost approaches. Only semi-autonomous vehicles in the control active intersection approaches are controlled for emission reduction.
The training environments are then partitioned into 28 partitions, considering the semi-autonomous vehicle penetration level, lane count, and traffic signal phase count as partition features. However, the environments in each partition may not form a uniform distribution. This may make environments sampled in consecutive training iterations differ significantly, requiring adaptive learning strategies. To overcome this, we adopt a synthetic environment distribution $\Phi_{T}$ derived from the full scenario set $\Phi$ for each partition with a uniform distribution for policy training as illustrated in Figure 2. Additional details about the training environment distributions can be found in Appendix 10.7.
Within partition MT-DRL: Training an MT-DRL policy per partition involves solving an eco-driving contextual Markov Decision Process (cMDP) where each scenario is an MDP within the cMDP. Furthermore, MDPs within a cMDP can take on either single-agent or multi-agent forms. We adopt multi-agent control, necessitating implicit or explicit coordination among controlled vehicles for better emission reductions. Consequently, each MDP in the cMDP is modeled as a Partially Observable Markov Decision Process (POMDP).
Eco-POMDP: A POMDP can be defined by extending an MDP when the agent’s knowledge is limited to observing a partial view of the actual state rather than directly perceiving the complete state. POMDP is characterized by a seven-tuple that expands the five-tuple MDP by incorporating the observation space and conditional observation probabilities. The definitions of observations, rewards, and actions of our POMDP formulation for each vehicle (agent) are as follows.
Observation: Vehicle observation design is mainly influenced by the currently available technology, such as the requisite vehicle sensors, and the significance attributed by existing literature to specific information required for eco-driving. Thus, we define the observation for a vehicle (agent) $i$ as follows.
  • speed of the ego-vehicle
  • relative distance to the traffic signal
  • traffic signal state (red, green, or yellow) for the current phase
  • time remaining in the current phase
  • time remaining until the traffic signal turns green for the second and third cycle
  • vehicle location (i.e., a flag indicating whether the vehicle is approaching the intersection, at the intersection, or exiting the intersection)
  • index of the vehicle’s current lane
  • vehicle’s intention to turn right, left, or go straight at the upcoming intersection
  • for both the follower and the leader vehicles on the same lane, adjacent right lane, and left lane of the ego-vehicle:
    • leader/follower speed
    • relative distance to the leader/follower
    • leader/follower turn signals status (turning right, left, or none)
Furthermore, we also append a context vector describing the underlying intersection to the observation defined above.
  • number of lanes
  • length of the ego-lane
  • signal timing plan for the traffic signal phase corresponding to the vehicle
  • speed limit of the ego-lane
  • penetration rate of semi-autonomous vehicles
Action: The longitudinal acceleration of the vehicle is controlled by the policy.
Reward: The reward function must encourage algorithmic generalization within the cMDP while considering various competing objectives like emission reduction and travel time efficiency. Thus, the reward function for each semi-autonomous vehicle is expressed as $r_{t}^{i}=v_{t}^{i}+\alpha\mathbbm{1}_{v^{i}_{t}<\tau}+\beta e_{t}^{i}$, where $v_{t}^{i}$ represents the velocity of vehicle $i$ at time $t$, and $e_{t}^{i}$ represents its CO2 emissions. In this equation, $\alpha$, $\beta$, and $\tau$ are hyperparameters. The indicator function $\mathbbm{1}_{v^{i}_{t}<\tau}$ indicates whether the vehicle is stopped, while the term $e_{t}^{i}$ encourages low emissions. The velocity term captures the effect on travel time.
It’s generally accepted that limiting accelerations and idling reduces emissions Mintsis et al. (2020). But this may create traffic overflows or reduce intersection throughput. To account for these observations, we devise two policies per partition ${\psi_{i}}$: one, denoted as $\pi_{\psi_{i}}^{1}$, penalize idling, while the other, $\pi_{\psi_{i}}^{2}$, places less weight on it. To achieve this we set $\alpha=35$, $\beta=3$ and $\tau=1$ for $\pi_{\psi_{i}}^{1}$ and $\alpha=10$, $\beta=3$ and $\tau=1$ for $\pi_{\psi_{i}}^{2}$.
Hierarchical controller architecture. While our MT-DRL policy controls vehicle accelerations, we leverage a rule-based default SUMO lane-changing controller for lane changes. It captures 1) strategic lane changes, 2) cooperative lane changes, 3) tactical lane changes, and 4) regulatory lane changes Erdmann (2014). At every time step, we first check if the lane-changing controller would predict a lane-changing maneuver for the vehicle. If a lane-changing maneuver is warranted, precedence is given to the lane-changing maneuvers, and the lane-changing controller takes over the control of the vehicle to control both longitudinal and lateral movements of the vehicle for that time step. Otherwise, we use the MT-DRL policy by default to control the longitudinal movements of the vehicle.
6.3 Impact Assessment
Impact assessment involves leveraging the trained eco-driving policies to assess the benefits of eco-driving in the representative scenario set $\Phi^{\prime}$. Let $\Pi$ be the trained policy set for all 28 partitions. We further employ the zero-shot generalizability of these policies to conduct the final impact assessment. In particular, for a given semi-autonomous vehicle penetration level, we have seven partitions by varying the lane count and phase count. Recall that we denote each partition as $\psi_{i}$ where $i\in\{1,\cdots,7\}$ when the semi-autonomous vehicle penetration level is given. Let $\Pi^{{}^{\prime}}=[(\pi_{\psi_{i}}^{1},\pi_{\psi_{i}}^{2})_{i\in\{1,\cdots,7\}}\cup\pi^{b}$] be the subset of trained policies for all seven partitions under a given semi-autonomous vehicle penetration level and the status quo driving baseline control policy $\pi^{b}$.
For a given scenario $\phi$ belonging to one of the seven clusters, we seek the best-performing policy $\pi^{*}_{\phi}$ as follows,
$\pi^{*}_{\phi}=\operatorname*{argmin}_{\pi^{k}\in\Pi^{{}^{\prime}}}(g(\pi^{k},\pi^{b},\phi))$ (15)
The function $g(\cdot)$ is defined as,
$g(\pi^{k},\pi^{b},\phi)=\begin{cases}f(\pi^{k},\phi)&\text{if network decomposition constraints are satisfied}\\ M&\text{otherwise}\end{cases}$ (16)
such that $f(\pi,\phi)$ denoting the function that captures the CO2 emission for the scenario $\phi$ under the control law $\pi$ and $M$ is large enough constant to indicate the policy $\pi^{k}$ is invalid due to any of the network decomposition constraint violations as listed in Equations 6, 7, and 8. If any of the constraints are not satisfied, we default to the status quo driving baseline control policy $\pi^{b}$. Below, we describe how we check for these constraint violations.
Control active intersection throughput constraint: To check if the control policy $\pi^{k}$ satisfies the throughput requirement, we compare its throughput with the status quo driving baseline $\pi^{b}$ throughput. If $n_{\pi^{k}}$ is the throughput of $\pi^{k}$ and $n_{\pi^{b}}$ is their throughput of $\pi^{b}$, then the constraint is satisfied if $n_{\pi^{k}}\geq n_{\pi^{b}}$.
Following intersection queue length to lane length ratio constraint: If the eco-driving increases the throughput of the intersections by smoothing traffic flow, it is imperative to ensure that additional throughput is bearable by the intersection immediately ahead of the control active intersection. To ensure this, we leverage the average vehicle queue length to lane length ratio of each approach as it indicates how many additional vehicles the approach can accommodate without congestion in the equilibrium state. In particular, we check if the vehicle inflow to all the control active intersections is increased by $r$ percent (mimicking increased throughput from the preceding intersection), can we ensure that all control active intersections under policy $\pi^{k}$ have vehicle queue length to lane length ratio $m_{\pi^{k}}\leq\tau$ where $\tau$ is a hyperparameter. When $\tau<1$, the lane can accommodate more vehicles. In practice, we find the maximum $r$ with $\tau=0.3$ such that at least 99% of the intersections can accommodate the additional throughput. Once such $r$ is found, all presented analyses will ensure that the following intersection queue length to lane length ratio constraint is satisfied.
Preceding intersection vehicle waiting time constraint: When eco-driving is performed at the control active intersection, the changed vehicle flow may affect the outflow of the intersection immediately precedes it (e.g., when the vehicles glide at the control active intersection slowing the rate other vehicles leave the preceding intersection). To ensure this does not happen, we check if $w_{\pi^{k}}\leq w_{\pi^{b}}$ where $w_{\pi^{k}}$ and $w_{\pi^{b}}$ are the average waiting time per vehicle in the preceding intersection under eco-driving policy $\pi^{k}$ and status quo driving baseline $\pi^{b}$, respectively. In simulations, this also includes any vehicle that has been scheduled to be spawned in the road network but yet to be spawned due to congested traffic settings.
Finally, by performing above zero-shot transfer scheme to all scenarios in $\Phi^{{}^{\prime}}$, we obtain the best policy $\pi^{*}_{\phi}$ for each scenario $\phi$. We then use the best policy of each scenario for subsequent impact assessments. In particular, regional eco-driving effectiveness $E(\cdot)$ is used to assess the overall benefit level as per Equation 4. Each scenario is simulated five times for statistical significance, and results are averaged following the recommended practice Agarwal et al. (2021).
7 Simulation-based Analysis
7.1 Experimental setup
We use Proximal Policy Optimization (PPO) Schulman et al. (2017) algorithm in training our MT-DRL policies with actor-critic architecture. The actor and critic do not share parameters and are separate neural networks. Centralized training and a decentralized execution paradigm are used for training. Section 10.8 in the Appendix describes the architecture details and all the hyperparameter configurations used in training.
7.2 Impact of eco-driving
Our eco-driving assessment encompasses diverse perspectives. First, we analyze the overall regional eco-driving effectiveness and the resultant variations in the intersection throughput in the three cities. Second, we analyze which intersection-related factors contribute greater influence at different levels of eco-driving adoption. Third, we analyze the level of intersection compatibility needed to achieve the benefits, informing implementation plans. Fourth, we analyze eco-driving behaviors and examine the safety implications of eco-driving, an aspect that has not received adequate attention in previous work Carslaw et al. (2010). Last, we contextualize eco-driving by analyzing its role in joint consideration with travel growth projections and decarbonization technologies such as vehicle electrification and the adoption of hybrid vehicles.
7.2.1 Eco-driving effectiveness
Figures 8 show the annual average effectiveness of eco-driving, assessing the percentage emission reduction and average throughput improvement. Eco-driving achieves a notable 11-22% emission reduction, equivalent to the national carbon emissions of Israel and Nigeria, respectively (refer to Appendix 10.2 for the estimation). Notably, a 10% eco-driving adoption can yield 25-50% of the total emission reduction, showcasing a non-linear benefit scaling and greater potential for short-term deployments to gain a majority of total benefits. Importantly, these gains incur no costs to intersection throughput. Additionally, Figure 9 demonstrates consistent emission benefits under different weather conditions across seasons, with minor fluctuations. This emphasizes that eco-driving can yield benefits throughout the year.
Figure 10 shows the spatial distributions of these emission benefits under different eco-driving adoption levels. The low benefits intersections in Figure 10 can happen due to three reasons. First, eco-driving at those intersections would not yield benefits due to their characteristics (e.g., not enough lane lengths for eco-driving). Second, the benefits of those intersections are defaulted to zero due to network decomposition constraint violation. Third, our MT-DRL policies are suboptimal in these intersections. Therefore, the benefits we observe in our analysis are a lower bound for the potential of eco-driving, and future work can address these challenges and increase the benefit levels.
Figure 8: Left Annual average emission reduction vs. eco-driving adoption (by factoring all seasons and peak and off-peak hours). Right Throughput improvement with varying adoption levels. By design, the throughput improvement is at least 0% when evaluating the emission reduction. Also, we ensure there throughput increases are not negatively affect the network level traffic.
Figure 9: Impact of weather conditions on eco-driving benefits. For each city, under every season, we present the eco-driving emission benefits with two different weather conditions: sunny and rainy. Not all cities go through every weather condition for a given season. We shift some of the curves within ±$\pm 2$% on the x-axis for better visibility.
7.2.2 Influential factors in eco-driving
Through a Pearson correlation analysis in Figure 11, we underscore key factors influencing emission benefits for informed intersection design and city planning. We consider nine key eco-driving factors that are known to have a significant impact on emission benefits and are directly observable, including temperature, humidity, speed limit, road grade, signal time ratio, vehicle inflow, phase count, lane count, and lane length. The signal ratio denotes the ratio between approach-related phase times and total cycle time. Notably, speed limits, vehicle inflow, and lane lengths significantly impact emission benefits. In particular, when the speed limit and lane length increase, the benefit levels increase. On the other hand, when the inflow increases, the benefit decreases. This explains why the three cities we consider have different benefit levels. As discussed in Appendix Section 10.5, different cities have different intersection feature distributions (refer to Figure 17 in the Appendix). Thus, cities like Atlanta demonstrate higher benefits, given their higher speed limits and longer incoming approaches, compared to relatively low-benefit cities like San Francisco.
Figure 10: Spatial distribution of eco-driving benefit at each of the three cities with varying eco-driving adoption levels.
Figure 11: Correlation analysis of eco-driving factors to the emission reduction benefits in all three cities under different eco-driving adoption levels.
Moreover, a noteworthy shift in influential factors is observed depending on the level of eco-driving adoption. In scenarios with lower adoption (10%), factors such as signal ratio and lane count carry more weight, whereas, at higher adoption levels (100%), speed limits, vehicle inflow, and lane length become more significant. This observation suggests that optimizing intersections for eco-driving is non-trivial, as benefits may either rise or decline with changing adoption rates over time. Furthermore, the impact of the signal ratio goes from a negative correlation to a positive correlation with the increasing adoption level. This highlights that traffic signal timing guidelines will need to be updated frequently over time with the adoption of eco-driving in order to retain the benefit levels.
7.2.3 Intersection compatibility
Figure 12 shows Pareto charts for all three cities revealing that at least 70% of emission benefits can be achieved by implementing eco-driving in just 20% of intersections at each adoption level. This emphasizes the potential for an efficient phased deployment, focusing on a specific subset of intersections. However, the Venn diagram in Figure 13 illustrates the specific 20% of intersections delivering the most significant emission benefits at every adoption level may have little overlap between eco-driving adoption levels. It highlights that intersections effective at lower eco-driving adoption levels may not maintain the same efficacy as eco-driving adoption increases, emphasizing the need for careful deployment planning, such as which intersections to be embedded with roadside units (RSU) for Signal and Phase Timing (SPaT) communication. In broad terms, this would question what infrastructure is considered long-term value for eco-driving as transportation agencies are often constrained on monetary budgets for infrastructure.
Figure 12: For each city, the four plots present Pareto charts displaying emission benefits vs intersections in each city at different eco-driving adoption levels. For each plot, the y-axis denotes the cumulative emission benefit percentage, and the x-axis denotes the percentage of intersections that produce a given cumulative benefit level. In plotting the x-axis, the intersections are ordered based on their emission reduction contributions.
Figure 13: The a,b, and c plots provide Venn diagrams for the three cities depicting the 20% of intersections that yield the most significant emission benefits under each eco-driving adoption level. These figures show how intersections evolve with changing adoption levels (the numbers on the figure are the intersection counts). A closer look at the figure underscores the need for meticulous planning when deploying eco-driving, as certain intersections that prove advantageous at lower eco-driving adoption rates may not exhibit the same level of effectiveness when adoption rates are high. Figures d,e, and f show the spatial distribution of the most effective 20% of intersections and how the overlaps between eco-driving adoption levels vary.
7.2.4 Eco-driving behaviors
In this section, we examine the behavioral changes in vehicles via learned policies compared to human-like driving. In Figure 14a, a two-lane representative intersection scenario with human-driven vehicles shows typical stopping at a traffic signal, resulting in unwanted idling. In contrast, policy $\pi^{1}$ in Figures 14b and 14d exhibits a non-stopping behavior, with most vehicles smoothly passing through the intersection without coming to a halt. Policy $\pi^{2}$ in Figure 14e mirrors this behavior but adopts a more human-like approach at lower adoption due to less weight on penalizing stopping, as seen in Figure 14c.
(a) 100% human-driven vehicles
(b) 10% eco-driving with $\pi^{1}$ policy
(c) 10% eco-driving with $\pi^{2}$ policy
(d) 100% eco-driving with $\pi^{1}$ policy
(e) 100% eco-driving with $\pi^{2}$ policy
Figure 14: Time-space diagrams of semi-autonomous vehicles (highlighted) and human-driven vehicles (muted) for a selected representative two-lane intersection.
We further analyze the induced road safety metrics of these learned behaviors compared to human-like driving. Figure 15 presents a roadway safety analysis using widely used surrogate safety measures (SSMs) commonly employed in assessing roadway safety Wang et al. (2021a). We analyze five SSMs, namely: time to collision (TTC), post-encroachment time (PET), maximum speed (MaxS), change in speed (DeltaS), and deceleration rate (DR). Further details on the analysis and used surrogate measures are in Appendix Section 10.9. Results in Figure 15 suggest that the learned behaviors are as safe as human-like driving, resulting from smoother vehicle behaviors. However, given that eco-driving behaviors of controlled vehicles are emerging behaviors that may come as counterintuitive for regular human drivers, it may trigger unforeseen human-driver responses, requiring additional human-factor studies for a better understanding of such human factors at play Jayawardana et al. (2023).
Figure 15: The safety of vehicle flow behavior measured using safety surrogate measures (SSM). The score for safety surrogate measures of learned policies in comparison to human-like driving. Each row is normalized by the human driving baseline $\pi^{b}$. Arrows indicate whether higher or lower is preferred.
Figure 16: Eco-driving benefits in the presence of vehicle electrification and hybrid engines as alternative decarbonization strategies. Figure shows total reductions over 2024-2050 in carbon emissions at various constant eco-driving adoption levels amidst projected electric vehicle and hybrid adoption in each city. Eco-driving yields additional benefits relative to a future with hybridization or electrification alone. For example, with an average total adoption rate of 20% from 2024 to 2050, eco-driving alone would cut San Francisco emissions in that period by about 7%, and when combined with projected eclectic vehicle and hybrid adoption, would cut emissions by about 17%. Results are relative to an all internal combustion engine vehicles baseline (i.e., without electrification or hybrid adoption.)
7.2.5 Role of eco-driving in the presence of alternatives
Next, we analyze the potential role of eco-driving in the presence of other popular transportation decarbonization strategies. Our analysis reveals a complementary relationship between eco-driving and the growing prevalence of hybrid and electric vehicle adoption. Figure 16 demonstrates the concurrent reduction in carbon emissions from increased eco-driving and the projected adoption of electric and hybrid vehicles in the US fleet up to the year 2050 for the three cities. Because the electric grid is not carbon-free, energy savings from eco-driving behavior continue to result in carbon emissions reductions even as more electric vehicles are adopted. Further details of the analysis can be found in Appendix Section 10.10.
8 Conclusion
Eco-driving has long captured the attention of the automotive industry, academics, city planners, and policymakers, yet a comprehensive understanding of its impact has remained elusive for decades, resulting in no large-scale deployments. The multitude of traffic scenario variations at scale has been a major obstacle, impeding progress in the field. In response to this longstanding call for a thorough and multi-factorial study that goes beyond existing efforts, we assess the prospective impact of eco-driving on a scale nearly four orders of magnitude larger than any previous study. Through our comprehensive analysis, we provide insightful perspectives on eco-driving, demonstrating its ability to yield substantial benefits in the pursuit of national climate change mitigation goals. We find the operational plans for eco-driving deployments require careful planning, accommodating the changing benefit dynamics with adoption rates. City planning and policy-making to support eco-driving are also informed through the analysis, especially in concert with other decarbonization strategies.
9 Acknowledgments
This research was partially supported by the Utah Department of Transportation (Project F-ST99(783)), and the authors extend their appreciation to Michael Sheffield, Christopher Siavrakas, and Kelly Njord at the Utah Department of Transportation for constructive discussions and providing their expert opinions on the work. Authors also extend their appreciation to Steven Barrett, Samitha Samaranayake, Zuduo Zheng, Yashar Farid, Jinhua Zhao, and Felice Frankel for their constructive comments. The authors also thank Sunera Avinash, Anirudh Valiveru, Jiaxin He, and Dajiang Suo for their help in numerous stages of the project. The authors acknowledge MIT SuperCloud and the Lincoln Laboratory Supercomputing Center for providing computational resources supporting the research results in this paper.